LUCIFER Commissioning (1a) and AGw Commissioning for 18-19 September 2008 (Day 21)
Observers: N. Ageorges, W. Seifert, and numerous other LUCIFERians (Peter, Kai, Marcus, Volker, Andreas) on mtn, D. Thompson.
J.Hill and D. Miller in Tucson
Software Support: N. Cushing, C. Biddick, M. De La Pena in Tucson
Telescope Operators: A. Ceranski
2:15 Opened at sunset
off-axis, Lucifer sees good round stars in direction of guide probe, but sausages in opposite side of the field. We are at 39 elevation. Could be secondary shift due to elevations causing binodal astigmatism.
collimate off-axis through wfsc 27
3:45 move on-axis and take 5 measurements /Repository/AGW_Data/080919w/ 1-5
moved on-axis to off-axis guide star, sent ~ 350 nm of focus: Should have been ~0
450 changed focus from 17.1 to 15.0, WFSC sent +300 nm focus, wrong direction
4:53 focus = 19. much smaller elongation on Lucifer
4:55 focus = 21. astigmatism small, but opposite axis: Went through focus.
5:00 set focus = 20.0 Collimate through 127
5:18 offset to star @ 90 degrees at about 4' to check if this is focal plane tilt. wfsc 128-151
=> conclusion is that either equation for focus correction as a function of off-axis distance is incorrect or we still have several hundred nm of focal plane tilt.
6:22 WFS loop failed to restart after pause: Found that Lucifer was sending offsetPointing every 40 seconds so 4 or 5 WFSC images were thrown away.
6:30 - 11:00 Lucifer march along with their commissioning plan
8:06 Power glitch. Aaron restarted mount, and restarted MCSPU. Killed MCS and PCS and restarted.
10:55 Network event: ECS died, the AGW lost communication with the PMAC. ECS restarted, stopAGW, startAGW, home all AGW axes. GCS hung. Killed GCS and restarted.
Serious focus errors. Aaron changed M1 Z global offset. We think active optics should have been cleared on both M1 and M2, but they were not, so a mirror collimate produced the bad focus.
12:15 Actuator 328 was noticed as disabled. We haven't searched the logs to see when it became disabled.
- 19 Sep 2008